# Late-blooming magnetars: awakening as long period transients after a dormant cooling epoch

### v2

@@ I Introduction
Radio surveys are steadily revealing a Galactic population of astrophysical bodies termed long period transients (LPTs) (Hurley-Walker et al., 2023; Caleb et al., 2024; Dong et al., 2025). Although a robust classification for this burgeoning class is yet to emerge, coherent radio pulsations from $\gtrsim$ 15 LPTs have been detected since the discovery of GLEAM-X J162759.5–523504 (henceforth GX J1627; Hurley-Walker et al., 2022). Optical spectroscopy has established that some LPTs are white dwarfs with type M companions (de Ruiter et al., 2025; Hurley-Walker et al., 2024). However, whether all LPTs contain dwarfs remains debated, as neutron stars – particularly magnetars – are compelling candidates for those that lack binary signatures (Suvorov & Melatos, 2023; Lyman et al., 2025) or exhibit interpulse (Lee et al., 2025) phenomena.
@@
The neutron-star hypothesis has garnered further support following the discovery of DART/ASKAP J1832–0911 (DA J1832; Li et al., 2024; Wang et al., 2025). This transient, with a period of $P\approx 44.2$ min, was detected simultaneously in the radio and X-ray bands by a host of instruments between December 2023 and September 2024, a discovery enabled by its serendipitous alignment with the supernova remnant G22.7–0.2. This marked the first instance of such multiwavelength activity – characteristic of transient magnetars that become radio-loud during outbursts (Turolla, Zane & Watts, 2015; Camilo et al., 2016) – in a LPT. The high radio luminosity (on-pulse $L_{\rm rad,max}\sim 10^{32}\text{ erg s}^{-1}$) and linear polarization ($\gtrsim 80\%$) of the source also disfavours a dwarf, as these features are difficult to explain with binary-instigated mechanisms (Melrose, 2017; Qu & Zhang, 2025).
@@
Aside from the X-ray burst observed from DA J1832, isolated LPTs share several features which have thus far eluded theoretical models. They are: (1) cold — there have been no persistent, soft X-ray detections to date; (2) slow — with periods spanning several minutes (Dong et al., 2025) to a few hours (Lee et al., 2025); and (3) irregularly active — often being radio-loud for only short windows within any given observational campaign (Hurley-Walker et al., 2022; Caleb et al., 2024). Reconciling all of these properties within a single framework has proven challenging: magnetars are categorically hot, young, and with spin periods $P\lesssim 12$ s (Mereghetti, Pons & Melatos, 2015; Coti Zelati et al., 2018).
@@
Magnetar phenomena are attributed to their strong magnetic fields (Thompson & Duncan, 1995, 1996). They are observationally defined by explosive eruptions in the form of bursts or flares, triggered by the redistribution of magnetic flux as the field evolves. Joule dissipation partially counteracts neutrino cooling, and advanced models incorporating this effect successfully reproduce the observed high luminosities (Viganò et al., 2013; Coti Zelati et al., 2018) and outburst rates (Göǧüş et al., 2000, 2001) as a function of age and magnetic intensity (see Pons & Viganò, 2019, for a review). The success of these magnetothermal models hinges on the proviso that electric currents circulate primarily within the crust of the star. This is because magnetic dissipation (and hence heating) is more effective there owing to the shorter length scales and presence of highly-resistive layers (Pons, Viganò & Rea, 2013). By contrast, simulations of neutron stars with fields of strength $\sim 10^{14}$ G and ‘core-threading’ (CT) currents exhibit essentially passive cooling (Dehman et al., 2023a) and predict little activity within the first $\gtrsim$ kyr of life (Dehman et al., 2020). Thus, observational considerations suggest that magnetars are born with magnetic configurations in which a substantial fraction of the sustaining electric currents reside in the solid crust, i.e. are predominantly crust-confined (CC). The CC/CT distinction should not be understood as a strict dichotomy, but rather as one end of a continuum of possible current distributions, ranging from configurations confined to the outer layers of the crust to those extending progressively deeper into the stellar interior. The apparent prevalence of CC configurations may be explained in two broad ways. One possibility is that magnetars are born with magnetic field configurations in which most of the sustaining currents are already concentrated in the crust. Alternatively, a substantial fraction of the magnetic flux may initially thread the core but be expelled to the crust at early times, rapidly establishing a crust-dominated configuration that can power magnetar activity on observable timescales (Bransgrove, Levin & Beloborodov, 2025; Lander et al., 2026).
@@
Magnetars typically slow down after outbursts (e.g., Archibald et al., 2017; Borghese et al., 2021; Lower et al., 2023). This implies that some fraction of the energy that powers bursts and flares also goes towards decreasing the rotational kinetic energy. Because magnetic decay restricts high-energy activity to early evolutionary stages in CC models however, magnetar spin periods are limited to at most $\sim 20$ s during their active life (Pons, Viganò & Rea, 2013).
@@ II.1 Crustal evolution
Within minutes after birth, nuclei in the outer layers of a neutron star freeze and form a solid crust (Yakovlev & Pethick, 2004; Potekhin, Pons & Page, 2015). Relativistic electrons within the crust carry a current that advects field lines through the Hall effect, generating small-scale structures that are especially susceptible to Ohmic dissipation (Mereghetti, Pons & Melatos, 2015; Dehman et al., 2023b). As hydrodynamical timescales are much shorter than either of these processes, the dynamics in the crust are well-described by the ‘electron-MHD’ induction equation,
@@
Only when $\mathcal{R}_{m}\gtrsim 10^{2}$ do Hall-related phenomena – where magnetic energy transfers between small and large scales – operate (Pons & Viganò, 2019). This, combined with the fact that both the thermal and electrical diffusivities depend on the temperature and field strength, necessitates a coupled magnetothermal evolution. We solve the relevant heat-diffusion equation together with (1), adopting the standard (Yakovlev & Pethick, 2004; Potekhin, Pons & Page, 2015; Dehman et al., 2023a) approach where the temperature at the crust-envelope boundary, $T_{b}$, is matched to a stationary envelope assuming radiative equilibrium. This links the thermal profile at the edge of the computational domain to the surface temperature, $T_{s}$, which determines the outgoing blackbody emission. The $T_{s}$-$T_{b}$ relation, which depends primarily on the average opacity and local gravity (Gudmundsson, Pethick & Epstein, 1983), serves as an outer boundary condition for $T(t,\boldsymbol{x})$. Technical details on the evolution scheme can be found in Appendix A.
Several evolutionary pathways can lead to neutron stars with magnetic fields that are significantly stronger in the crust (Torres-Forné et al., 2016; Barrère et al., 2022; Bransgrove, Levin & Beloborodov, 2025) or even absent entirely (Lander, 2024; Lander et al., 2026) from the core. For such stars (i.e., those called CC in this paper), core-field evolution is irrelevant, and the above equations fully describe the system given a background stellar model. Throughout this paper, we adopt the BSk24 equation of state (Pearson et al., 2018) for a ‘canonical’ star with a mass of $M=1.4M_{\odot}$, radius $R=12.426$ km, and crustal thickness of $0.86$ km.
II.2 Core evolution
If the core field is strong enough to be dynamically important, one must also specify how $\boldsymbol{B}$ and $T$ evolve there. The core composition is unknown, but relative motions between charged particles and neutrals define an ambipolar velocity which transports flux in a manner analogous (Goldreich & Reisenegger, 1992; Viganò et al., 2021) to the Hall effect. Because these streaming motions are inherently hydrodynamical however, it is unclear which terms (if any) can be safely discarded in the multi-fluid equations describing momentum conservation. Stemming largely from the application of different approximation schema – as modelling the full system over astrophysical timescales is impossible – there are disagreements in the literature about the relevance of ambipolar diffusion (see Pons & Viganò, 2019, for a review). Even in ‘optimistic’ scenarios (e.g. Moraga et al., 2024) though, significant drifts likely take longer than a Myr for large-scale fields of strength $\sim 10^{14}$ G unless the core is very cold and the direct Urca process operates (Passamonti et al., 2017).
We neglect ambipolar drift and evolve the magnetic and thermal fields of stars with ‘moderate’ fields over Myr timescales. In CT models, the core field is formally evolved using the same induction equation as in the crust to ensure a smooth match at the crust–core interface. However, the effective Reynolds number in the core is very small and Hall dynamics are suppressed (Goldreich & Reisenegger, 1992), so the core field remains essentially frozen over the timescales considered. Therefore, the core evolution has a negligible impact on the crustal dynamics and observable quantities, and could equivalently be replaced by a fixed-core boundary condition. Our approach is employed only for numerical convenience: it allows for a smooth matching of the magnetic field across the crust–core interface and avoids potential artifacts at the boundary. The currents are essentially chained to the core for CT models.
@@ III Outbursts and late-time awakening
The main driving idea is that, during a Hall-dominated epoch, flux waves of large amplitude can drift into malleable regions of the outer crust. The otherwise-rigid ion lattice is sheared as a result, and a mechanical strain is built up (Thompson & Duncan, 1995; Thompson, Lyutikov & Kulkarni, 2002). If a critical strain threshold is breached, a crustquake or some type of failure can occur that liberates magnetoelastic energy (e.g., Lander & Gourgouliatos, 2019; Dehman et al., 2020). By tracking stresses throughout a given simulation we identify the failure sites. In particular, consider the (traceless) Maxwell stress tensor,
@@
for local stress limit $\sigma^{\rm max}$. These stresses are tracked throughout the simulations with event properties recorded when the inequality (4) is met at any given spatial location and time step ($\sim 0.1$ yr). Note that, in many previous studies (Perna & Pons, 2011; Pons & Perna, 2011; Dehman et al., 2020), the full stress tensor was used rather than the traceless component (3). The latter is arguably more appropriate for predicting failures via deviatoric stresses (Lander & Gourgouliatos, 2019) from a continuum mechanics perspective; the adjusted criterion used here is more conservative as magnetic pressures cannot directly contribute to overstraining (the trace of the full tensor is twice the magnetic pressure). A quantitative comparison between several different failure criteria is provided in Appendix A.1.
To determine the energetics of an event, we integrate the square of the relative difference between the magnetic field at the failure time and the reference state, defined as the field at the start of the current equilibrium epoch (i.e., defining $M_{ij}^{\rm eq}$ just after the previous failure). The critical stress threshold for crustal failure remains somewhat uncertain, as does the actual failure mechanism itself. If the crust behaves as a body-centred crystal though, molecular dynamics (Chugunov & Horowitz, 2010) simulations indicate that
@@
for Coulomb parameter $\Gamma$ (electrostatic-to-thermal energy ratio), atomic charge $Z$, ion-number density $n_{i}$, and where $a$ is the typical inter-nuclear distance ($\propto n_{i}^{-1/3}$). The affected region readjusts to a new equilibrium, resolidifies, and then resumes its elastic response, accumulating stress until another rupture (Perna & Pons, 2011; Pons & Perna, 2011). Failures can avalanche as adjacent regions exceeding a threshold,
@@
become unstable at sites determined by the crests of Hall or thermoplastic waves (Beloborodov & Levin, 2014). In previous works (e.g., Dehman et al., 2020), values in the range $0.8\lesssim\epsilon\lesssim 1$ are typically used; we fix $\epsilon=0.9$.
@@ III.1 Radio activity cycle
Although the exact nature of failures is poorly understood, because of the immense hydrostatic pressures it is thought that, rather than shattering like glass as brittle material would, failed regions of the crust deform plastically (Beloborodov & Levin, 2014; Lander & Gourgouliatos, 2019). Fluid circulating within ‘plastic islands’ surrounding failure sites constitutes an effective longitudinal velocity which, via induction, produces an electric field that works to accelerate charges to high Lorentz factors. This initiates pair cascading and can trigger radio activity for relatively modest fields ($\gtrsim 10^{14}$ G; Cooper & Wadiasingh, 2024). One can think of the plastic flow as a localised circulation which dominates over rigid rotation for LPTs. In particular, radio-activation theory demands a minimum voltage drop ($\Delta V\propto BP^{-2}$) which is difficult to achieve for LPTs unless $B\gg 10^{14}$ G because $P$ is so large (Chen & Ruderman, 1993; Suvorov & Melatos, 2023). Therefore, the star can only become radio-loud when a failure occurs and a plastic flow is initiated (see Cooper & Wadiasingh, 2024). The relative ‘activity window’ set by the duration of a plastic flow and frequency of failures defines a physical duty cycle, $D^{\rm p}$, which can be quantified.
For CT cases – the proposed LPT track – the waiting time distribution peaks at $\sim 1$ yr, and thus if motions circulate in a plastic island for $\sim 6$ months (as anticipated from the radio continuity of DA J1832 and other LPTs), we estimate an upper limit of $D^{\rm p}_{\rm max}\sim 0.5$. On the other hand, given the narrowness of plastic islands, an observer will catch the beam only after events that are favourably positioned with respect to their line of sight. Even if the physical duty cycle is large therefore – in the sense that a cone of radiation is emitted perpendicular to the plastic island that forms after each failure – the observed duty cycle, $D$, is reduced by a solid angle (Cooper & Wadiasingh, 2024). For example, if the star is positioned such that only failures occurring at a narrow range of co-latitudes surrounding $\theta\sim 0.3$ rad are visible to a given observer, the right panel of Fig. 3 indicates that only 1 in $\gtrsim 10^{2}$ events may reveal in radio (i.e., $D\sim 0.5\%$). Such a configuration would explain the observed (Dong et al., 2025) value of $D\sim 0.6\%$ from CHIME J0630+25 and other isolated LPTs which display duty cycles of order $\sim 1\%$ (see Table 1 and Appendix B). While it is difficult to falsify the model based on $D$ values because LPT viewing geometries are unconstrained, one prediction of the model is clear: X- and radio-band activity should go together since both originate from the same event.

### v3

@@ I Introduction
Radio surveys are steadily revealing a Galactic population of astrophysical bodies termed long period transients (LPTs) (Hurley-Walker et al., 2023; Caleb et al., 2024; Dong et al., 2025). Although a robust classification for this burgeoning class is yet to emerge, coherent radio pulsations from $\gtrsim$ 15 LPTs have been detected since the discovery of GLEAM-X J162759.5–523504 (henceforth GX J1627; Hurley-Walker et al., 2022). Optical spectroscopy has established that some LPTs are white dwarfs with type M companions (de Ruiter et al., 2025; Hurley-Walker et al., 2024). However, whether all LPTs contain dwarfs remains debated, as neutron stars – particularly magnetars – are compelling candidates for those that lack binary signatures (Suvorov and Melatos, 2023; Lyman et al., 2025) or exhibit interpulse (Lee et al., 2025) phenomena.
@@
The neutron-star hypothesis has garnered further support following the discovery of DART/ASKAP J1832–0911 (DA J1832; Li et al., 2024; Wang et al., 2025). This transient, with a period of $P\approx 44.2$ min, was detected simultaneously in the radio and X-ray bands by a host of instruments between December 2023 and September 2024, a discovery enabled by its serendipitous alignment with the supernova remnant G22.7–0.2. This marked the first instance of such multiwavelength activity – characteristic of transient magnetars that become radio-loud during outbursts (Turolla et al., 2015; Camilo et al., 2016) – in a LPT. The high radio luminosity (on-pulse $L_{\rm rad,max}\sim 10^{32}\text{ erg s}^{-1}$) and linear polarization ($\gtrsim 80\%$) of the source also disfavours a dwarf, as these features are difficult to explain with binary-instigated mechanisms (Melrose, 2017; Qu and Zhang, 2025).
@@
Aside from the X-ray burst observed from DA J1832, isolated LPTs share several features which have thus far eluded theoretical models. They are: (1) cold — there have been no persistent, soft X-ray detections to date; (2) slow — with periods spanning several minutes (Dong et al., 2025) to a few hours (Lee et al., 2025); and (3) irregularly active — often being radio-loud for only short windows within any given observational campaign (Hurley-Walker et al., 2022; Caleb et al., 2024). Reconciling all of these properties within a single framework has proven challenging: magnetars are categorically hot, young, and with spin periods $P\lesssim 12$ s (Mereghetti et al., 2015; Coti Zelati et al., 2018).
@@
Magnetar phenomena are attributed to their strong magnetic fields (Thompson and Duncan, 1995, 1996). They are observationally defined by explosive eruptions in the form of bursts or flares, triggered by the redistribution of magnetic flux as the field evolves. Joule dissipation partially counteracts neutrino cooling, and advanced models incorporating this effect successfully reproduce the observed high luminosities (Viganò et al., 2013; Coti Zelati et al., 2018) and outburst rates (Göǧüş et al., 2000, 2001) as a function of age and magnetic intensity (see Pons et al., 2025, for a review). The success of these magnetothermal models hinges on the proviso that electric currents circulate primarily within the crust of the star. This is because magnetic dissipation (and hence heating) is more effective there owing to the shorter length scales and presence of highly-resistive layers (Pons et al., 2013). By contrast, simulations of neutron stars with fields of strength $\sim 10^{14}$ G and ‘core-threading’ (CT) currents exhibit essentially passive cooling (Dehman et al., 2023a) and predict little activity within the first $\gtrsim$ kyr of life (Dehman et al., 2020). Thus, observational considerations suggest that magnetars are born with magnetic configurations in which a substantial fraction of the sustaining electric currents reside in the solid crust, i.e. are predominantly crust-confined (CC). The CC/CT distinction should not be understood as a strict dichotomy, but rather as one end of a continuum of possible current distributions, ranging from configurations confined to the outer layers of the crust to those extending progressively deeper into the stellar interior. The apparent prevalence of CC configurations may be explained in two broad ways. One possibility is that magnetars are born with magnetic field configurations in which most of the sustaining currents are already concentrated in the crust [^1] A scenario explaining magnetar formation for a CC field configuration that accounts for the chiral magnetic effect is presented in Dehman and Pons (2025).. Alternatively, a substantial fraction of the magnetic flux may initially thread the core but be expelled to the crust at early times, rapidly establishing a crust-dominated configuration that can power magnetar activity on observable timescales (Bransgrove et al., 2025; Lander et al., 2026).
@@
Magnetars typically slow down after outbursts (e.g., Archibald et al., 2017; Borghese et al., 2021; Lower et al., 2023). This implies that some fraction of the energy that powers bursts and flares also goes towards decreasing the rotational kinetic energy. Because magnetic decay restricts high-energy activity to early evolutionary stages in CC models however, magnetar spin periods are limited to at most $\sim 20$ s during their active life (Pons et al., 2013).
@@ II.1 Crustal evolution
Within minutes after birth, nuclei in the outer layers of a neutron star freeze and form a solid crust (Yakovlev and Pethick, 2004; Potekhin et al., 2015). Relativistic electrons within the crust carry a current that advects field lines through the Hall effect, generating small-scale structures that are especially susceptible to Ohmic dissipation (Mereghetti et al., 2015; Dehman et al., 2023c, b). As hydrodynamical timescales are much shorter than either of these processes, the dynamics in the crust are well-described by the ‘electron-MHD’ induction equation,
@@
Only when $\mathcal{R}_{m}\gtrsim 10^{2}$ do Hall-related phenomena – where magnetic energy transfers between small and large scales – operate (Pons et al., 2025; Dehman and Brandenburg, 2025). This, combined with the fact that both the thermal and electrical diffusivities depend on the temperature and field strength, necessitates a coupled magnetothermal evolution. We solve the relevant heat-diffusion equation together with (1), adopting the standard (Yakovlev and Pethick, 2004; Potekhin et al., 2015; Dehman et al., 2023a) approach where the temperature at the crust-envelope boundary, $T_{b}$, is matched to a stationary envelope assuming radiative equilibrium. This links the thermal profile at the edge of the computational domain to the surface temperature, $T_{s}$, which determines the outgoing blackbody emission. The $T_{s}$-$T_{b}$ relation, which depends primarily on the average opacity and local gravity (Gudmundsson et al., 1983), serves as an outer boundary condition for $T(t,\boldsymbol{x})$. Technical details on the evolution scheme can be found in Appendix A.
Several evolutionary pathways can lead to neutron stars with magnetic fields that are significantly stronger in the crust (Torres-Forné et al., 2016; Barrère et al., 2022; Bransgrove et al., 2025) or even absent entirely (Lander, 2024; Lander et al., 2026) from the core. For such stars (i.e., those called CC in this paper), core-field evolution is irrelevant, and the above equations fully describe the system given a background stellar model. Throughout this paper, we adopt the BSk24 equation of state (Pearson et al., 2018) for a ‘canonical’ star with a mass of $M=1.4M_{\odot}$, radius $R=12.426$ km, and crustal thickness of $0.86$ km.
II.2 Core evolution
@@
If the core field is strong enough to be dynamically important, one must also specify how $\boldsymbol{B}$ and $T$ evolve there. The core composition is unknown, but relative motions between charged particles and neutrals define an ambipolar velocity which transports flux in a manner analogous (Goldreich and Reisenegger, 1992; Viganò et al., 2021) to the Hall effect. Because these streaming motions are inherently hydrodynamical however, it is unclear which terms (if any) can be safely discarded in the multi-fluid equations describing momentum conservation. Stemming largely from the application of different approximation schema – as modelling the full system over astrophysical timescales is impossible – there are disagreements in the literature about the relevance of ambipolar diffusion (see Pons et al., 2025, for a review). Even in ‘optimistic’ scenarios (e.g. Moraga et al., 2024) though, significant drifts likely take longer than a Myr for large-scale fields of strength $\sim 10^{14}$ G unless the core is very cold and the direct Urca process operates (Passamonti et al., 2017).
We neglect ambipolar drift and evolve the magnetic and thermal fields of stars with ‘moderate’ fields over Myr timescales. In CT models, the core field is formally evolved using the same induction equation as in the crust to ensure a smooth match at the crust–core interface. However, the effective Reynolds number in the core is very small and Hall dynamics are suppressed (Goldreich and Reisenegger, 1992), so the core field remains essentially frozen over the timescales considered. Therefore, the core evolution has a negligible impact on the crustal dynamics and observable quantities, and could equivalently be replaced by a fixed-core boundary condition. Our approach is employed only for numerical convenience: it allows for a smooth matching of the magnetic field across the crust–core interface and avoids potential artifacts at the boundary. The currents are essentially chained to the core for CT models.
@@ III Outbursts and late-time awakening
The main driving idea is that, during a Hall-dominated epoch, flux waves of large amplitude can drift into malleable regions of the outer crust. The otherwise-rigid ion lattice is sheared as a result, and a mechanical strain is built up (Thompson and Duncan, 1995; Thompson et al., 2002). If a critical strain threshold is breached, a crustquake or some type of failure can occur that liberates magnetoelastic energy (e.g., Lander and Gourgouliatos, 2019; Dehman et al., 2020). By tracking stresses throughout a given simulation we identify the failure sites. In particular, consider the (traceless) Maxwell stress tensor,
@@
for local stress limit $\sigma^{\rm max}$. These stresses are tracked throughout the simulations with event properties recorded when the inequality (4) is met at any given spatial location and time step ($\sim 0.1$ yr). Note that, in many previous studies (Perna and Pons, 2011; Pons and Perna, 2011; Dehman et al., 2020), the full stress tensor was used rather than the traceless component (3). The latter is arguably more appropriate for predicting failures via deviatoric stresses (Lander and Gourgouliatos, 2019) from a continuum mechanics perspective; the adjusted criterion used here is more conservative as magnetic pressures cannot directly contribute to overstraining (the trace of the full tensor is twice the magnetic pressure). A quantitative comparison between several different failure criteria is provided in Appendix A.1.
To determine the energetics of an event, we integrate the square of the relative difference between the magnetic field at the failure time and the reference state, defined as the field at the start of the current equilibrium epoch (i.e., defining $M_{ij}^{\rm eq}$ just after the previous failure). The critical stress threshold for crustal failure remains somewhat uncertain, as does the actual failure mechanism itself. If the crust behaves as a body-centred crystal though, molecular dynamics (Chugunov and Horowitz, 2010) simulations indicate that
@@
for Coulomb parameter $\Gamma$ (electrostatic-to-thermal energy ratio), atomic charge $Z$, ion-number density $n_{i}$, and where $a$ is the typical inter-nuclear distance ($\propto n_{i}^{-1/3}$). The affected region readjusts to a new equilibrium, resolidifies, and then resumes its elastic response, accumulating stress until another rupture (Perna and Pons, 2011; Pons and Perna, 2011). Failures can avalanche as adjacent regions exceeding a threshold,
@@
become unstable at sites determined by the crests of Hall or thermoplastic waves (Beloborodov and Levin, 2014). In previous works (e.g., Dehman et al., 2020), values in the range $0.8\lesssim\epsilon\lesssim 1$ are typically used; we fix $\epsilon=0.9$.
@@ III.1 Radio activity cycle
Although the exact nature of failures is poorly understood, because of the immense hydrostatic pressures it is thought that, rather than shattering like glass as brittle material would, failed regions of the crust deform plastically (Beloborodov and Levin, 2014; Lander and Gourgouliatos, 2019). Fluid circulating within ‘plastic islands’ surrounding failure sites constitutes an effective longitudinal velocity which, via induction, produces an electric field that works to accelerate charges to high Lorentz factors. This initiates pair cascading and can trigger radio activity for relatively modest fields ($\gtrsim 10^{14}$ G; Cooper and Wadiasingh, 2024). One can think of the plastic flow as a localised circulation which dominates over rigid rotation for LPTs. In particular, radio-activation theory demands a minimum voltage drop ($\Delta V\propto BP^{-2}$) which is difficult to achieve for LPTs unless $B\gg 10^{14}$ G because $P$ is so large (Chen and Ruderman, 1993; Suvorov and Melatos, 2023). Therefore, the star can only become radio-loud when a failure occurs and a plastic flow is initiated (see Cooper and Wadiasingh, 2024). The relative ‘activity window’ set by the duration of a plastic flow and frequency of failures defines a physical duty cycle, $D^{\rm p}$, which can be quantified.
For CT cases – the proposed LPT track – the waiting time distribution peaks at $\sim 1$ yr, and thus if motions circulate in a plastic island for $\sim 6$ months (as anticipated from the radio continuity of DA J1832 and other LPTs), we estimate an upper limit of $D^{\rm p}_{\rm max}\sim 0.5$. On the other hand, given the narrowness of plastic islands, an observer will catch the beam only after events that are favourably positioned with respect to their line of sight. Even if the physical duty cycle is large therefore – in the sense that a cone of radiation is emitted perpendicular to the plastic island that forms after each failure – the observed duty cycle, $D$, is reduced by a solid angle (Cooper and Wadiasingh, 2024). For example, if the star is positioned such that only failures occurring at a narrow range of co-latitudes surrounding $\theta\sim 0.3$ rad are visible to a given observer, the right panel of Fig. 3 indicates that only 1 in $\gtrsim 10^{2}$ events may reveal in radio (i.e., $D\sim 0.5\%$). Such a configuration would explain the observed (Dong et al., 2025) value of $D\sim 0.6\%$ from CHIME J0630+25 and other isolated LPTs which display duty cycles of order $\sim 1\%$ (see Table 1 and Appendix B). While it is difficult to falsify the model based on $D$ values because LPT viewing geometries are unconstrained, one prediction of the model is clear: X- and radio-band activity should go together since both originate from the same event.